# MedHalu: Hallucinations in Responses to Healthcare Queries by Large Language Models

### v1

@@ Abstract
The remarkable capabilities of large language models (LLMs) in language understanding and generation have not rendered them immune to hallucinations. LLMs can still generate plausible-sounding but factually incorrect or fabricated information. As LLM-empowered chatbots become popular, laypeople may frequently ask health-related queries and risk falling victim to these LLM hallucinations, resulting in various societal and healthcare implications.
In this work, we conduct a pioneering study of hallucinations in LLM-generated responses to real-world healthcare queries from patients. We propose MedHalu, a carefully crafted first-of-its-kind medical hallucination dataset with a diverse range of health-related topics and the corresponding hallucinated responses from LLMs with labeled hallucination types and hallucinated text spans. We also introduce MedHaluDetect framework to evaluate capabilities of various LLMs in detecting hallucinations. We also employ three groups of evaluators – medical experts, LLMs, and laypeople – to study who are more vulnerable to these medical hallucinations.
We find that LLMs are much worse than the experts. They also perform no better than laypeople and even worse in few cases in detecting hallucinations. To fill this gap, we propose expert-in-the-loop approach to improve hallucination detection through LLMs by infusing expert reasoning. We observe significant performance gains for all the LLMs with an average macro-F1 improvement of 6.3 percentage points for GPT-4.
@@ 1 Introduction
Large Language Models (LLMs) have made significant strides towards artificial general intelligence, achieving notable success in healthcare [1, 2, 3], finance [4, 5, 6], and law [7] , exemplified by models like GPT-3.5/4 [8], Gemini [9], LLaMA [10], and Pixtral [11]. Despite these advancements, LLMs often suffer from hallucination, producing factually incorrect information that is deceptive, nonsensical, or unfaithful to the source content, raising safety concerns and hindering their deployment [12, 13]. As LLM-powered chatbots like ChatGPT [8] gain prominence among the general public, laypeople with no healthcare background increasingly seek health-related advice from these models. This unconditional trust makes them vulnerable to the hallucinated information generated by LLMs.
@@
Existing works in LLM hallucinations in medical domains [14, 15, 16] mainly focus on testing the medical knowledge of LLMs through standardized medical exam questions, which may not sufficiently demonstrate how these models interact with end users in the real world for the following reasons: 1) contextual dependency — Interactions with real-world users often involve ambiguous or incomplete queries, requiring the model to interpret the missing context, which increases the risk of generating hallucinated information. In contrast, medical exams typically present well-defined, clear-cut questions with definitive answers. 2) data diversity — Real-world queries come from end users with diverse backgrounds and varying levels of understanding. This is different, for example, from LLMs answering medical exams, which contain standardized questions and answers that do not reflect the variability found in real-world interactions. Considering this diversity and ambiguity, it is unclear how accurate and reliable answers can be provided by LLMs and experts alike.
This Work. For the first time in the literature, we study hallucinations in LLM generated responses to real-world healthcare queries. To address the twin issues of contextual dependency and data diversity, we first gather a dataset consisting of extensive real-world queries from users with a diverse set of backgrounds, ranging from experts (the HealthQA dataset [17]) to laypeople (LiveQA [18] and MedicationQA [19]). The questions exhibit characteristics such as 1) ambiguity: open-ended or vague queries that require inference of additional context; 2) incomplete information: questions lacking necessary details; 3) user diversity: queries from users with varying levels of medical knowledge and different ways of expressing their concerns. Inspired by [20], we categorize and adapt three types of hallucinations – input-conflicting, context-conflicting and fact-conflicting hallucinations specifically for healthcare question-answer pairs. We then create our dataset, MedHalu, which consists of hallucinations containing LLM-generated answers to healthcare queries. To create these answers, we write prompts that ask LLMs explicitly to generate hallucinated responses based on the three hallucination types. Each answer in MedHalu is labeled with the hallucination types and the hallucinated text spans for each of the LLM generated responses.
@@
We then validate the answers among native English speakers, both laypeople and experts (having at least an undergraduate degree in health or medicine), asking each to judge whether a given LLM-generated answer contains a hallucination. We find that, as expected, laypeople (average macro-F1 score of ) are less likely to be able to detect hallucinations than experts (average macro-F1 of ).
The difficulty that laypeople experience in detecting hallucinations raises obvious safety concerns about the inevitable use of LLMs for healthcare queries. As a first attempt at mitigating such concerns, we evaluate whether LLMs can themselves be trained to detect hallucinations. We experiment with various LLMs, including both open-source (Llama-2) and proprietary models (GPT-3.5 and GPT-4), to study their capabilities of detecting hallucinations using prompting. Surprisingly, we find that LLMs are able to detect hallucinations only at a similar level to laypeople and significantly underperform compared to human experts (e.g., GPT-4 only gets an average macro-F1 score of ).
We therefore go back to the human experts, and ask them to articulate their reasoning about how they detected hallucinations. Their responses indicated that a combination of domain knowledge as well as looking up trusted health-related public websites such as UpToDate, BMJBestPractice, WebMD and NHS (UK) were used to cross-verify the answers. Based on this, we propose an expert-in-the-loop approach to infuse expert reasoning into the prompts for LLMs to detect hallucinations, and show that this greatly increases the performance of automatically detecting hallucinations using LLMs.
Contribution. Our contributions are as follows:
@@
Dataset. We introduce MedHalu, the first-of-its-kind benchmark specifically designed to study LLM hallucinations in real-world healthcare queries. MedHalu features question-answer pairs from diverse health topics together with fine-grained hallucination types and the hallucinated text spans.
@@
Method. We propose MedHaluDetect, a hallucination detection framework, and conduct comprehensive experiments with both open-source and proprietary LLMs (e.g., GPT-3.5/4) to measure their hallucination detection capabilities using prompts.
@@
Empirical Findings. We conduct a holistic comparison of the capabilities and vulnerabilities in hallucination detection across three groups of evaluators – LLMs, medical experts, and laypeople. Our findings reveal that LLMs perform no better than laypeople in detecting hallucinations. In contrast, medical experts excel at identifying medical hallucinations and significantly outperform LLMs. To fill this gap, we propose expert-in-the-loop approach to improve medical hallucination detection through LLMs by infusing expert reasoning into the prompt. This method results in notable performance improvements across all LLMs, with an average macro-F1 increase of 6.3% for GPT-4.
We aim to facilitate ongoing research and development in the field by releasing all of our code, data, and tools upon the acceptance of this work. The paper is structured as follows: Section 2 discusses LLM hallucination definition, hallucination types and examples. Section 3 introduces the proposed MedHalu benchmark and the dataset statistics. Section 4 introduces MedHaluDetect framework for medical hallucination detection and discusses the results. In section 5, we introduce expert-in-the-loop approach to improve the capabilities of LLMs in detecting medical hallucinations. Finally, we discuss related works in Section 6 and conclusions and future directions in Section 7.
2 LLM Hallucinations in Healthcare Queries
We focus on LLM hallucinations in response to healthcare queries from real-world patients. Despite the impressive text generation capabilities of LLMs, they often encounter challenges in generating factually accurate content. Hallucinations, as defined by [21], refer to generated content that is nonsensical or unfaithful to the provided source content. Inspired by [20], we categorize and adapt the following types of hallucinations specifically for healthcare question-answer pairs.
Fact-conflicting Hallucination: When the LLM generated answer to the healthcare query conflicts with a well-known fact or universal truth.
@@ 2 LLM Hallucinations in Healthcare Queries
Input-conflicting Hallucination: When the LLM generated answer conflicts with the healthcare query asked (deviates from the source input provided by users).
Context-conflicting Hallucination: When the LLM generated answer conflicts with its own generated text (self-conflict).
An example for each hallucination type is shown in Table 1. Sample healthcare queries are taken from HealthQA and LiveQA datasets. Hallucinated answers are generated using OpenAI’s GPT-3.5 [22] model. Corresponding hallucinated text spans are highlighted in red color. Response to the first question is an example of input-conflicting hallucination since it fails to answer the question appropriately. Second response to the question regarding foods that cause most allergies is a context-conflicting hallucination. It is because LLM first generates that fresh fruits and vegetables can cause allergies and then towards the end says that a diet of fresh fruits can minimize the risk of allergic reactions. Therefore, the LLM is conflicting with its own answer (self-conflict) in this example. Similarly, third response to the causes of pneumonia is an example of fact-conflicting hallucination since the answer contains some phrases that are factually incorrect (e.g., pneumonia is caused by exposure to cold temperatures).
3 MedHalu Dataset
In this section, we introduce the first LLM-generated MedHalu benchmark, its generation methodology (Section 3.1), human evaluation to assess the quality (Section 3.2) and the dataset statistics (Section 3.3).
@@ 3.1 Dataset Generation
| I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect. You should write the hallucinated answer using the following method: <hallucination type definition>. You are trying to answer a question but there is a <hallucination type> hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example: <An example healthcare query, expert answer and the hallucinated answer.> You should try your best to make a hallucinated answer to the following question: Question: <Healthcare query> Correct Answer: <Expert answer> Hallucinated Answer: |
Our novel MedHalu benchmark is designed to study LLM hallucinations in real-world healthcare queries based on the three prominent publicly-available healthcare datasets, each containing healthcare-related questions and answers curated by medical experts.
LiveQA [18] contains question-answer pairs from real consumer health questions received by the U.S. National Library of Medicine (NLM).
MedicationQA [19] contains anonymous consumer questions, primarily related to drugs and medication from MedlinePlus. The answers are sourced from trusted websites such as MedlinePlus, DailyMed, Mayo Clinic, etc.
These datasets were selected for their diverse health topics and expert-verified answers, aligning with real-world public healthcare queries.
For each question in each dataset, we used GPT-3.5 [8] to generate hallucinated answers to each healthcare query. We design specific prompts for each type of hallucination and derive hallucinated answers from the GPT-3.5 model. Table 2 shows the template of hallucination generation prompt. For individual prompts for each hallucination type, please refer to Appendix A. For each query and the hallucination type (i.e., the three types mentioned in Section 2), the corresponding prompt from Table 2, along with the healthcare query, was input into GPT-3.5 to generate hallucinated answers, which were then labeled with the hallucination type.
@@ 3.2 Human Evaluation
To ensure that the LLM-generated healthcare responses are hallucinated with the specific hallucination types specified in the prompt during the generation process, we conduct human annotations. We employ six medical experts through Prolific111https://www.prolific.com, last accessed 8 Aug 2024. with an hourly rate of US. Medical experts are selected only if they are native English speakers and have graduated with at least an undergraduate degree in health or medicine. To follow our institute’s ethics guidelines, all the annotations took place in the United Kingdom.
To keep the costs of annotations in check, we randomly sample around of MedHalu dataset ( question-answer pairs) using stratified sampling of the base datasets – HealthQA, LiveQA, and MedicationQA. We randomly split question-answer pairs into batches, each containing pairs. We then hire experts for each batch to evaluate question-answer pairs in two hours. Specifically, we designed and deployed an annotation platform based on our specific needs in our institute’s research server. We provide each annotator with detailed annotation guidelines and a video tutorial of the annotation platform and ask their consent to share basic information such as their education background. We then ask each annotator to annotate every question-answer pair. They are asked whether the answer contains a hallucination and if yes, the type of hallucination following the definitions provided in the guidelines. We also introduce attention checks randomly to a subset of question-answer pairs and have only accepted annotations from those experts who passed all the attention checks.
On average, expert annotators achieve a Cohen’s Kappa score of which denotes substantial agreement of the LLM-generated responses with the expert responses. High agreement denotes that GPT-3.5 has indeed generated healthcare responses pertaining to specific hallucination types and therefore, the MedHalu dataset is reliable along with hallucination type labels.
3.3 Dataset Statistics
| Hallucination Type | HQA | LQA | MQA | Total |
@@ 3.3 Dataset Statistics
Table 3 shows the statistics of MedHalu dataset for each of the hallucination types. “None-conflict” signifies that the healthcare answer does not contain any hallucination. MedHalu dataset contains question-answer pairs in total, out of which pairs are from HealthQA, are from LiveQA and pairs are from MedicationQA. The dataset is balanced with roughly similar number of question-answer pairs in each of the four hallucination types for all the three base datasets.
@@ 4 Hallucination Detection in Healthcare Queries
Detecting LLM hallucinations is a challenging task mainly because the generated content may seem to be plausible and semantically similar to the correct answer. In this section, we discuss our hallucination detection framework – MedHaluDetect (Section 4.1), experimental setup (Section 4.2), and results (Sections 4.3 and 4.4).
@@ 4.1 Methodology
To detect LLM hallucinations in healthcare queries, we introduce our MedHaluDetect framework. At first, we employ three groups of evaluators: LLMs, medical experts, and laypeople without any healthcare background to detect the presence of hallucinations in MedHalu question-answer pairs. By comparing their evaluations, we study the extent to which each group is vulnerable to hallucinated healthcare responses. With the release of ChatGPT and other LLM-powered chatbots to the general public, laypeople often ask health-related queries from these models, trust them without any healthcare background, and therefore become vulnerable to hallucinated information.
@@
Given a healthcare query and corresponding response from MedHalu, we ask these three groups of evaluators whether any type of hallucination is involved in the response and the reason behind their decisions. If hallucinations are detected, we further ask these evaluators to highlight specific text spans where these hallucinations occur to assess the granularity of their detection abilities.
For hallucination detection using LLMs, we prompt various LLMs to identify the presence of hallucinations and specific hallucinated text spans based on the definitions of different hallucination types, the healthcare query, and the corresponding response from the MedHalu dataset. The prompt for hallucination detection is detailed in Table 4. We employ the following models for detecting LLM hallucinations: the open-source LLaMA-2-7B [10] model and two proprietary GPT-3.5 [22] and GPT-4 [23] models.
We employ groups of medical experts and laypeople through Prolific. Medical experts are selected only if they are native English speakers and have graduated with at least an undergraduate degree in health/medicine. Only those laypeople are selected who are also native English speakers but do not have any degree or background in healthcare/medicine. In order to keep the costs of human evaluation in check, we randomly sampled the MedHalu dataset using stratified sampling of the base datasets – HealthQA, LiveQA, and MedicationQA. We sample question answer pairs in total. We randomly split question-answer pairs into batches, each containing pairs. We hire evaluators for each batch to evaluate question-answer pairs in two hours. Therefore, we employ six medical experts and six laypeople in the overall study. Please refer to Section 3.2 for detailed annotation process.
4.2 Experimental Setup and Evaluation Metrics
Experimental Setup. For generating hallucinated responses to the healthcare queries, we use GPT-3.5 [22] using OpenAI’s official API. We set temperature to and maximum generation length to tokens. For detecting LLM hallucinations, we input our detection prompt into each of the LLMs together with the healthcare query and corresponding response. For LLaMA-2-Instruct [10], we use its open-source implementation after downloading the weights for model with billion parameters. For OpenAI’s GPT-3.5 and GPT-4, we use their official API. We set the same temperature of and maximum generation length of for all the LLMs.
Evaluation Metrics. We model hallucination detection as a binary classification task and therefore, we use accuracy, macro-Precision, macro-Recall and macro-F1 scores to evaluate the performance across different LLMs and different evaluators. We also ask these evaluators to highlight hallucinated text spans. To measure the effectiveness of detecting hallucinated text spans, we use edit distance between the LLM-detected text spans and the expert annotated hallucinated spans. Edit distance is a measure of similarity between the two strings and is defined as the minimum number of changes required to convert first string into another by inserting, deleting or replacing the characters. The smaller the edit distance, the more similar the strings are.
4.3 Results
@@
| LLM | Mean | Median | Mean | Median | Mean | Median |
@@
Table 5 shows the results for different groups of evaluators for hallucination detection on MedHalu dataset. For HealthQA subset, LLaMA-2 achieves an F1-score of whereas GPT-3 and GPT-4 obtains F1-scores of and respectively. For LiveQA subset, the highest F1-score is only for both GPT-3.5 and GPT-4. This is because LiveQA is a challenging dataset since it contains real consumer health queries received by the U.S. National Library of Medicine. For MedicationQA, the highest F1-score is . Overall, OpenAI’s GPT-3.5 and GPT-4 perform much better than the open source LLaMA-2 model for hallucination detection with only a slight improvement in GPT-4 over GPT-3.5 model.
Table 5 also shows the aggregated results for groups of medical experts and laypeople employed for hallucination detection. Medical experts get an overall macro-F1 score of for HealthQA, only for LiveQA, and for MedicationQA. Note that low accuracy and macro-F1 scores even for the experts consistently suggest that hallucination detection in LiveQA dataset is challenging. Laypeople without any medical background only get macro-F1 scores of , and for HealthQA, LiveQA and MedicationQA subsets, respectively. As expected, laypeople perform much worse than the experts and therefore, are more vulnerable to these hallucinated healthcare responses. Surprisingly, LLMs perform no better than laypeople except in the LiveQA subset. Moreover, for HealthQA and MedicationQA, LLMs perform even worse than laypeople. This shows that LLMs are nowhere near the experts in medical hallucination detection due to the lack of domain knowledge required for this challenging task. Moreover, they are incapable of detecting self-generated hallucinated responses to the healthcare queries.
We next evaluate the capabilities of different LLMs for detecting hallucinated text spans. We consider the expert annotated hallucinated text spans for question-answer pairs as ground truth. We then calculate the edit distance between all the possible combinations of LLM-detected and expert annotated text spans and select the minimum score for each text span detected by an LLM. Table 6 shows the mean and median edit distance values between the LLM-detected and the expert-detected hallucinated text spans for each of the subsets in MedHalu dataset. We exclude LLaMA-2 because it was incapable of detecting hallucinated text spans during our initial experiments even though we tried with various different prompts. GPT-3.5 achieves mean edit distance values of , , and for HealthQA, LiveQA and MedicationQA, respectively. On the other hand, GPT-4 achieves mean edit distance values of , , and for HealthQA, LiveQA and MedicationQA, respectively. Out of GPT-3.5 and GPT-4 models, GPT-4 consistently has a higher agreement with expert evaluators as evident from its lower edit distance values. LiveQA gets the highest edit distance values among the three subsets for both GPT-3.5 and GPT-4, again indicating that LiveQA is a challenging subset. For the benefit of the research community, we will also make these LLM detected hallucinated text spans publicly available to allow fine-grained hallucination detection.
4.4 Hallucination Detection per Hallucination Type
| LLM | Hallucination
Type |
HealthQA | LiveQA | MedicationQA | |||
| Accuracy | macro-F1 | Accuracy | macro-F1 | Accuracy | macro-F1 | ||

### v2

@@ Abstract
Large language models (LLMs) are starting to complement traditional information seeking mechanisms such as web search. LLM-powered chatbots like ChatGPT are gaining prominence among the general public. AI chatbots are also increasingly producing content on social media platforms. However, LLMs are also prone to hallucinations, generating plausible yet factually incorrect or fabricated information. This becomes a critical problem when laypeople start seeking information about sensitive issues such as healthcare. Existing works in LLM hallucinations in the medical domain mainly focus on testing the medical knowledge of LLMs through standardized medical exam questions which are often well-defined and clear-cut with definitive answers. However, these approaches may not fully capture how these LLMs perform during real-world interactions with patients.
This work conducts a pioneering study on hallucinations in LLM-generated responses to real-world healthcare queries from patients. We introduce MedHalu, a novel medical hallucination benchmark featuring diverse health-related topics and hallucinated responses from LLMs, with detailed annotation of the hallucination types and text spans. We also propose MedHaluDetect, a comprehensive framework for evaluating LLMs’ abilities to detect hallucinations. Furthermore, we study the vulnerability to medical hallucinations among three groups — medical experts, LLMs, and laypeople. Notably, LLMs significantly underperform human experts and, in some cases, even laypeople in detecting medical hallucinations. To improve hallucination detection, we propose an expert-in-the-loop approach that integrates expert reasoning into LLM inputs, significantly improving hallucination detection for all LLMs, including a macro-F1 improvement for GPT-4.
@@ 1 Introduction
Large Language Models (LLMs) have made significant strides towards artificial general intelligence, achieving notable success in domains such as healthcare (Cascella et al. 2023; Xu et al. 2024; Chen et al. 2024c), finance (Wu et al. 2023), and law (Cui et al. 2023), exemplified by models like GPT-4 (Achiam et al. 2023), GPT-3.5 (Ouyang et al. 2022), and LLaMA (Touvron et al. 2023). Despite these advancements, LLMs often suffer from hallucination, producing factually incorrect information that is deceptive, nonsensical, or unfaithful to the source content, raising safety concerns and hindering their deployment (Rawte, Sheth, and Das 2023).
For at least 10 years now, it has been recognised that people seek out answers to highly sensitive and personal medical information on the web and social media (De Choudhury, Morris, and White 2014). Researchers are starting to recognise that LLMs can complement traditional information seeking mechanisms like web search (Fernández-Pichel, Pichel, and Losada 2024). Indeed, as LLM answers start to get incorporated into Google search, the distinction between search and LLMs is starting to get blurred.
As LLM-powered chatbots like ChatGPT (OpenAI 2023a) gain prominence among the general public, laypeople with no healthcare background increasingly seek health-related advice from these models (Ayers et al. 2023). Even educated adults find it difficult to tell whether some information is generated by AI chatbots (KFF 2024). This unconditional trust makes them vulnerable to the hallucinated information generated by LLMs.
@@
In this rapidly evolving information landscape, it becomes important to guard against the known pitfalls of LLMs such as hallucinations. Systematically measuring the dangers of hallucination in the real world is a hard problem. Existing works in LLM hallucinations in the medical domain (Pal, Umapathi, and Sankarasubbu 2023; Chen et al. 2024a; Vishwanath et al. 2024) focus mainly on testing the medical knowledge of LLMs through standardized medical exam questions. These approaches may not fully capture how these models perform in real-world interactions because of two important reasons: 1) Contextual Dependency. Real-world user queries can be ambiguous or incomplete, requiring the models to infer missing context, which increases the risk of generating hallucinated information. 2) User Variability. Real-world queries originate from end users with diverse backgrounds and varying levels of medical understanding. This is significantly different from LLMs answering standardized medical exams, which contain standardized well-defined, clear-cut questions with definitive answers that do not reflect the variability in real-world interactions. Thus, the accuracy and reliability of LLM responses in real-world scenarios remain an open question.
@@
This Work. We present the first study of LLM hallucinations in responses to real-world healthcare queries. To address the dual challenges of contextual dependency and user variability, we first collect a dataset of extensive real-world queries from users with varying medical expertise, ranging from healthcare professionals (HealthQA (Zhu et al. 2019)) to laypeople (LiveQA (Abacha et al. 2017) and MedicationQA (Abacha et al. 2019)). The questions feature: 1) ambiguity, where queries are open-ended or vague with multiple meanings, requiring inference of additional context; 2) incomplete information, where essential details are missing; and 3) user diversity, reflecting varying levels of medical knowledge.
Inspired by Zhang et al. (2023), we categorize hallucinations into three types—input-conflicting, context-conflicting, and fact-conflicting—for healthcare question-answer pairs. We then create MedHalu, a medical hallucination detection benchmark consisting of hallucinated LLM-generated answers to healthcare queries. Each answer in MedHalu is labeled with the hallucination type and the corresponding hallucinated text spans.
Insights. Our findings reveal a significant gap in hallucination detection between laypeople (macro-F1: ) and experts (macro-F1: ), highlighting safety concerns about relying on LLMs for healthcare advice.
We then ask whether LLMs can be recursively used to detect hallucinations. For example, if a layperson obtains an answer from an LLM to a sensitive healthcare query, we ask LLM, if that generated answer is provided as input to (the same or a different) LLM, whether this provided piece of text contains hallucinations.
We find that unfortunately LLMs are no better than laypeople. Indeed, they perform at a similar level (macro-F1: for GPT-4) to laypeople and significantly underperform human experts at detecting text containing hallucinations.
We then dig deeper and ask human experts what strategies they use to identify hallucination-containing text. We find that human experts rely on a combination of domain knowledge and trusted health resources such as UpToDate111https://www.wolterskluwer.com/en/solutions/uptodate, last accessed 9 Jan 2025., BMJBestPractice222https://bestpractice.bmj.com/info/us/, last accessed 9 Jan 2025., WebMD333https://www.webmd.com/, last accessed 9 Jan 2025., and NHS444https://www.nhs.uk/, last accessed 9 Jan 2025. to cross-validate the answers. Building on this, we propose an expert-in-the-loop approach that integrates expert reasoning into LLM prompts, significantly improving LLMs’ ability to automatically detect hallucinations.
I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect.
You should write the hallucinated answer using the following method: <hallucination type definition>. You are trying to answer a question but there is a <hallucination type> hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example: <An example healthcare query, expert answer and the hallucinated answer.> You should try your best to make a hallucinated answer to the following question: Question: <Healthcare query> Correct Answer: <Expert answer> Hallucinated Answer: |
Contribution. Our key contributions are:
@@
Novel Dataset. We introduce MedHalu, the first medical hallucination detection benchmark specifically designed to study LLM hallucinations in real-world healthcare queries, featuring question-answer pairs from diverse health topics as well as fine-grained hallucination types and text spans.
@@
Comprehensive Framework. We propose MedHaluDetect, a hallucination detection framework, and conduct evaluation across both open-source models and proprietary LLMs (e.g., GPT-3.5/4) to measure their detection capabilities.
@@
Empirical Findings. We conduct a holistic comparison of the capabilities and vulnerabilities in hallucination detection across three groups of evaluators—LLMs, medical experts, and laypeople. Our findings reveal that LLMs perform no better than laypeople in detecting hallucinations. In contrast, medical experts excel at identifying medical hallucinations and significantly outperform LLMs.
Mitigation Strategy. To address this gap, we propose an expert-in-the-loop approach that integrates expert reasoning into LLM prompts, enhancing hallucination detection and resulting in improvements across all models and an average macro-F1 increase of % for GPT-4.
2 The MedHalu Benchmark
MedHalu is designed to study LLM hallucinations in real-world healthcare queries that are often asked by patients and laypeople on web forums. This section details the types of hallucinations we test for (Section 2.1), dataset generation process (Section 2.2), and human evaluation (Section 2.3).
2.1 Hallucination Types
Hallucinations occur when LLMs generate content that is nonsensical or unfaithful to the input (Ji et al. 2023). Inspired by Zhang et al. (2023), we identify three types of hallucinations. Examples for each type are shown in Table 1.
Fact-conflicting: the response contradicts a well-known fact or universal truth.
@@ 2.1 Hallucination Types
Input-conflicting: the response conflicts with or deviates from the input query.
Context-conflicting: the response is self-contradictory or internally inconsistent.
@@ 2.2 Dataset Generation
MedHalu is a medical hallucination detection benchmark based on the three publicly available, expert-curated healthcare datasets described below. Each dataset consists of open-ended healthcare-related questions and answers curated by medical experts.
HealthQA (Zhu et al. 2019) contains healthcare question-answer pairs constructed from healthcare articles on the popular health-services website Patient555https://patient.info, last accessed 9 Jan 2025.. The questions are created by medical experts from a diverse set of healthcare topics, and answers are sourced from the articles. It has an average (± standard deviation) query length of 7.72 ± 2.41 words. An example query is “What are the symptoms of tonsillitis?”.
LiveQA (Abacha et al. 2017) contains question-answer pairs from real consumer health questions received by the U.S. National Library of Medicine (NLM) with an average query length of 41.76 ± 37.38 words. An example query is “I work at the Airport and I am in contact with hundreds of people in a day. I want to be sure that I am not going to expose people to Shingles. When is it safe for me to go back to work?”.
MedicationQA (Abacha et al. 2019) contains anonymous consumer questions, primarily related to drugs and medication from MedlinePlus666https://medlineplus.gov, last accessed 9 Jan 2025. with an average query length of 6.86 ± 2.83 words. The answers are sourced from trusted websites such as MedlinePlus, DailyMed, Mayo Clinic, etc. An example query is “How many mg does it take to overdose on Oxycodone?”.
We select these three different subsets to have healthcare queries with diverse topics for hallucination dataset since the three datasets differ in terms of how these queries were obtained, the nature and the length of these queries. These queries align with real-world public healthcare queries. For each such healthcare query, we explicitly ask an LLM to generate hallucinations. Specifically, we design tailored prompts for each hallucination type (See Section 2.1) and feed them to GPT-3.5 (OpenAI 2023a) to capture answers that are expected to contain hallucinations.
Details of the hallucination generation prompts can be found in Table 2 and Appendix A. By carefully constructing these prompts, the MedHalu dataset captures a wide range of plausible healthcare-related hallucinations. This approach ensures that the generated examples are both realistic and relevant. While it may not encompass every possible real-world variation, we believe this provides an important starting point for assessing and mitigating hallucinations in the healthcare domain. We fully expect that future work will build on and improve on these, perhaps by adding additional kinds of queries or additional types of hallucinations.
@@ 2.3 Human Evaluation
To validate hallucinations in the LLM-generated responses, we employ medical experts through Prolific777https://www.prolific.com, last accessed 9 Jan 2025. with an hourly rate of US$18. Selected annotators were required to be native English speakers based in the UK with at least an undergraduate degree in health or medicine.
| Background | No. of Experts | |
| Education | Undergrad in Medicine | 14 |
| Graduate in Medicine | 16 | |
| Gender | Male | 18 |
| Female | 12 | |
| Other | 0 | |
| Race | White | 12 |
| Asian | 11 | |
| Black | 5 | |
| Other | 2 | |
To keep a balance between the cost and the quality, we randomly sample about of MedHalu ( question-answer pairs) using stratified sampling from HealthQA, LiveQA, and MedicationQA. These were split into batches of pairs each, with three medical experts assigned to each batch to conduct the evaluation in two hours. We designed and deployed a custom annotation platform (details in Appendix B), provided detailed annotation guidelines and a video tutorial, and obtained consent from the annotators to collect basic information about their background. Table 3 provides details of these medical experts and their background. This study also received approval from the ethics committee of our institution. For each question-answer pair, the experts were asked whether the answer contains any hallucination and the type of hallucination following the definitions provided in the guidelines. Experts were also asked to mark the hallucinated text spans in the answers for fine-grained hallucination detection. We also implemented random attention checks, and only the experts passing all these checks have their annotations accepted. The expert annotators achieve an average Cohen’s Kappa score of , denoting almost perfect agreement between experts and LLM-generated responses. This confirms the reliability of MedHalu as the LLM has indeed generated healthcare responses pertaining to specific hallucination types. The statistics of MedHalu for each hallucination type is shown in Table 4.
| Hallucination | HQA | LQA | MQA | Total |
@@ 3 Hallucination Detection in Healthcare Queries
Detecting LLM hallucinations is particularly challenging because the generated content may seem to be plausible and semantically similar to the correct answer. In this section, we discuss our hallucination detection framework—MedHaluDetect (Section 3.1), experimental setup (Section 3.2), and evaluation metrics (Section 3.3).
@@ 3.1 Methodology
Our MedHaluDetect framework for detecting LLM hallucinations in healthcare queries leverages input from three groups of evaluators: LLMs, medical experts, and laypeople without healthcare expertise. Given a healthcare query and its corresponding response from MedHalu, each group assesses whether the response contains any types of hallucination and provides justifications for their decisions. When hallucinations are detected, we further ask these evaluators to highlight specific text spans where these hallucinations occur to assess the granularity of their detection abilities.
Hallucination detection using LLMs. We prompt various models to identify the presence of hallucinations and the corresponding text spans based on the definitions of different hallucination types, the healthcare query, and the corresponding response from MedHalu. We employ both open-source models such as LLaMA-2 (Touvron et al. 2023) and proprietary models like GPT-3.5 (Ouyang et al. 2022) and GPT-4 (Achiam et al. 2023). The prompt for hallucination detection is detailed in Table 5. By comparing evaluations from all groups, we study their varying susceptibility to hallucinated healthcare responses.
@@
| LLM | Mean | Med | Mean | Med | Mean | Med |
@@
| HealthQA | LiveQA | MedicationQA | |||||
| LLM | Hallucination | Acc | ma-F1 | Acc | ma-F1 | Acc | ma-F1 |